OpenStudy (anonymous):

How many solutions can be found for the equation 4x − 8 − 4x = 9? Zero One Two Infinitely many

OpenStudy (anonymous):

How many solutions can be found for the equation 4x − 8 − 4x = 9? Zero One Two Infinitely many

OpenStudy (anonymous):

idk why its doing that ill type it out

Directrix (directrix):

Solve the equation for x: 4x − 8 − 4x = 9

OpenStudy (anonymous):

4x - 8 - 4x = 9

Directrix (directrix):

What is 4x -4x ?

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

is that what it means? zero answers possible?

Directrix (directrix):

4x − 8 − 4x = 9 4x − 4x - 8 = 9 0 - 8 = 9 ---> What next?

Directrix (directrix):

On the day that -8 =9, the equation will have a solution. That day will not be soon.

Directrix (directrix):

>>is that what it means? zero answers possible? YES

OpenStudy (anonymous):

so its impossible, ok then

OpenStudy (anonymous):

thanks

Directrix (directrix):

You are welcome.